OpenStudy (anonymous):

solve for x. show all steps leading to the solution 5x-2x+x=15

OpenStudy (anonymous):

x(5-2+1)=15 x(4)=15 x=15/4

OpenStudy (anonymous):

5x -2x +x = 15 (5-2+1)x = 15 4x =15 x = 15/4

OpenStudy (amistre64):

add, then divide